OpenStudy (anonymous):

What is the equation of a horizontal line passing through the point (2, 10)?

OpenStudy (anonymous):

In which form? Standard? Slope-intercept?

OpenStudy (anonymous):

standard form

OpenStudy (anonymous):

y = 12 x = −8 y = 10 x = 2

OpenStudy (anonymous):

those are the answer choices, but i dont understand how to do it

OpenStudy (anonymous):

We know that all horizontal lines have a slope of 0, so we can find our answer from there! We will start with the point-slope form on an equation. y-10 = 0 (x-2) Then we multiply.... y-10=0. Subtract y=10. Standard form is ax+by=c (btw) Because the line is horizontal you really do not have to go through all these steps, if you were to graph it, you would see that the x-coordinates do not really matter when it comes to answering this question. Anywho, the answer is y=10 :) Hope this helps!

OpenStudy (anonymous):

ahhh, thanks! that helped alot for me to understand it better now, thanks again :)